### v2

# Social Bias in Popular Question-Answering Benchmarks

Abstract
Question-answering (QA) and reading comprehension (RC) benchmarks are essential for assessing the capabilities of large language models (LLMs) in retrieving and reproducing knowledge. However, we demonstrate that popular QA and RC benchmarks are biased and do not cover questions about different demographics or regions in a representative way, potentially due to a lack of diversity of those involved in their creation. We perform a qualitative content analysis of 30 benchmark papers and a quantitative analysis of 20 respective benchmark datasets to learn (1) who is involved in the benchmark creation, (2) how social bias is addressed or prevented, and (3) whether the demographics of the creators and annotators correspond to particular biases in the content. Most analyzed benchmark papers provided insufficient information regarding the stakeholders involved in benchmark creation, particularly the annotators. Notably, just one of the benchmark papers explicitly reported measures taken to address social representation issues. Moreover, the data analysis revealed gender, religion, and geographic biases across a wide range of encyclopedic, commonsense, and scholarly benchmarks. More transparent and bias-aware QA and RC benchmark creation practices are needed to facilitate better scrutiny and incentivize the development of fairer LLMs.
Social Bias in Popular Question-Answering Benchmarks
Angelie Kraft University of Hamburg Leuphana University Lüneburg Weizenbaum Institute angelie.kraft@leuphana.de                        Judith Simon University of Hamburg                        Sonja Schimmler TU Berlin Fraunhofer FOKUS Weizenbaum Institute
1 Introduction
Large language models (LLMs) inhabit the core of a wide range of user-facing systems, powering applications such as chatbots, which are utilized as writing and coding assistants, search engines, and advisors. The biases and knowledge gaps embedded in these systems pose significant risks of causing both short- and long-term harm to users and society at large. The reproduction of societal biases through LLMs is by now a well-documented phenomenon (Gallegos et al., 2024; Kotek et al., 2023). Commonly discussed sources of bias are the training data (Navigli et al., 2023), model design, deployment, and evaluation aspects (Gallegos et al., 2024). Indeed, optimizing LLMs to perform well on popular benchmarks is highly incentivized, as strong performance can enhance a researcher’s visibility and credibility (Koch et al., 2021). However, if these widely used benchmarks are biased, they effectively incentivize model optimization towards biased standards (Bowman and Dahl, 2021). Our work provides one of the first systematic analyses demonstrating that many of the most widely adopted LLM benchmarks are, in fact, quietly unrepresentative.
Raji et al. (2021, p. 2) "describe a benchmark as a particular combination of a dataset or sets of datasets […], and a metric, conceptualized as representing one or more specific tasks or sets of abilities." They argue that popular benchmarks – while claiming universality – are actually limited in coverage and validity. Citing Haraway (2016), they criticize that this universality claim masks an inevitable positionality and value-ladenness which manifests itself in the lack of coverage of "non-Western contexts," under-representation of non-cis gender identities, and non-white racial identities. Yet, comprehensive empirical analyses of benchmark biases are sparse and mostly limited to benchmarks that are themselves dedicated to the measurement of bias instead of downstream task performance (Powers et al., 2024; Demchak et al., 2024).
Our work aims to fill this gap and focuses on question-answering (QA) and reading comprehension (RC) benchmarks, i.e., tasks where the model is presented an explicit question and its generated answer is then checked for correctness (e.g., open-ended, fill-in-the-gap, or multiple choice;  Rogers et al., 2023). We argue that these tasks are the most direct proxies to the ways in which users query chatbots to gather information and, thus, the ways in which LLMs are shaping modern knowledge ecosystems.
Building on the definition of an AI benchmark introduced by Raji et al. (2021), we define a socially biased QA or RC benchmark as one that exhibits a statistical skew in the occurrence of demographic and/or geographic identifiers or names within its dataset, corresponding to societal gradients of power and injustice, such as the under-representation of non-cis-male gender identities or non-Western individuals, locations, or events. We argue that such biases in QA and RC benchmarks can cause societal harm. By overlooking marginalized demographics and geographies in evaluation, these benchmarks encourage the optimization of knowledge-driven language technologies to favor the interests of a privileged few, exacerbating injustices related to knowledge, i.e., epistemic injustice Fricker (2007). This injustice unfolds in two forms: hermeneutic injustice, where marginalized groups lack the knowledge resources to interpret their social experiences, and testimonial injustice, where their credibility is diminished due to prejudiced judgments (Fricker, 2007). Biased benchmarks not only widen gaps in collective knowledge resources but also privilege certain knowledges over others, designating them as more desirable for LLMs to reproduce.
Our work seeks to answer the following research questions:
  • RQ1
    Who is involved in the creation of popular QA and RC benchmarks?
  • RQ2
    Are potential social biases avoided or addressed in the creation of the benchmarks?
  • RQ3
    Are potential social biases in the datasets reflected in the demographics of the individuals involved in the benchmark creation process?
Based on a manual analysis of the 30 most popular QA and RC benchmark papers as well as a quantitative data analysis of 20 benchmark datasets, we identified (a) a lack of transparency regarding the individuals involved in benchmark dataset creation, (b) a lack of intentional prevention of biases, and (c) prominent gender, occupation, religion, and location biases for encyclopedic, commonsense, maths, and science benchmarks. In unveiling these issues, our work adds to the mounting criticism of current AI evaluation practices and shines a light on biased benchmarks being a potential source of LLM bias by incentivizing biased inference heuristics. [^1] The source code can be found here: https://github.com/krangelie/qa-benchmark-biases
2 Related Works
LLMs reproduce stereotypical associations (Nadeem et al., 2021; Kotek et al., 2023) and achieve different levels of accuracy for examples referring to different social groups in downstream-tasks (Park et al., 2018; Kiritchenko and Mohammad, 2018), such as QA (Parrish et al., 2022; Jin et al., 2024). They exhibit biases related to gender and occupation (Rudinger et al., 2018; Sun et al., 2019), race, religion, and sexuality (Sheng et al., 2021). These biases can lead to representational and allocational harms Barocas et al. (2017); Blodgett et al. (2020). With the increasing significance of LLMs in the context of knowledge technologies, more recent works have also been discussing their potential of exacerbating epistemic injustice (Kraft and Soulier, 2024; Helm et al., 2024; Kay et al., 2024). Sources of bias are non-representative training data, the training or inference algorithm, the deployment context and user interface, as well as evaluation with unrepresentative benchmarks (Gallegos et al., 2024; Suresh and Guttag, 2021; Bowman and Dahl, 2021). Age, gender, race, educational background, and first language of an annotator can influence their annotations and, consequently, the ground truths used to train and evaluate models (Pei and Jurgens, 2023; Al Kuwatly et al., 2020). Hence, crowdworker samples with low demographic diversity produce datasets of correspondingly low diversity and generalizability (Geva et al., 2019). Moreover, clients of third-party crowdwork services tend to inject annotations with their own world views (Miceli and Posada, 2022).
Transparent documentation practices of datasets, including their biases and limitations, have been promoted as an important measure to prevent harmful outcomes (Bender and Friedman, 2018; Stoyanovich and Howe, 2019; Gebru et al., 2021), i.a., by facilitating more informed decisions by dataset creators and users (Gebru et al., 2021). Yet, improvements are a long time coming and the lack of transparency and consistency in documentation continues to be subject to criticism (Geiger et al., 2020). Reuel et al. (2024), recently proposed a structured AI benchmark assessment with criteria addressing aspects of design, implementation, documentation, maintenance, and retirement. They applied this schema to 24 foundation and non-foundation model benchmarks, covering natural language processing (NLP), agentic and ethical behavior benchmarks and found overall low levels of reproducibility and interpretability. MMLU scores lowest in their overall assessment. Our work sits in the same category but targets a social bias-related appraisal of benchmarks. Researchers in the area of algorithmic bias research have been investigating the biases of bias benchmarks, like BBQ (Powers et al., 2024; Parrish et al., 2022), BOLD and SAGED (Demchak et al., 2024; Dhamala et al., 2021; Guan et al., 2024). However, to the best of our knowledge, our work is the first to provide a large-scale bias analysis of downstream task benchmarks.
3 Method
Figure 1: Qualitative content analysis process for the benchmark papers.
3.1 Benchmark Selection
To identify popular QA and RC benchmarks, we firstly selected all benchmarks including textual data (not excluding multimodal datasets) in the Papers with Code (PwC) corpus of machine learning dataset metadata [^2] https://paperswithcode.com/about, accessed: September 17, 2024 and ranked them by their citation counts. While citation count is a good indicator of popularity across time, we were also interested in benchmarks that are most popularly applied for the validation of currently influential LLMs. To identify such, we selected the most highly ranked models on the Chatbot Arena LLM Leaderboard (Chiang et al., 2024), [^3] https://lmarena.ai/, accessed: September 18, 2024 as well as the language models with the most likes on HuggingFace. [^4] https://huggingface.co/models?sort=likes, accessed: September 13, 2024 We extracted the top 20 models from both lists and collected all of the 40 related reports, i.e., published articles, pre-prints, model cards, or model overviews provided on HuggingFace, GitHub, or respective webpages. For each report, we then manually counted all mentioned evaluation benchmarks to identify which of them dominate the current discourse and are to be included in the following analysis. Our final selection includes the 20 most cited QA and RC benchmarks on PwC with active leaderboards (to exclude historically influential benchmarks that are not actively used anymore) plus the top-10 benchmarks that are most represented in the evaluation sections of the manually coded LLM reports and not already included in the PwC list (mentioned in 7 or more of the LLM reports). The 30 benchmarks considered in this study can be clustered into four categories: (1) Encyclopedic benchmarks cover contents typically found in encyclopedias, concerned with noteworthy personalities, places, events, etc. Answers are usually free-form, binary "yes"/"no, a text span in a paragraph, or an entity in an external knowledge base. (2) Commonsense benchmarks pose questions about everyday knowledge, e.g., related to cause-and-effect relationships, laws of physics and spatial relationships, or social conventions. Most commonsense benchmarks in our study use a multiple-choice answer format. (3) Scholarly benchmarks are single- or multi-domain, based on academic exams or curricula, openly accessible educational resources, or authored by students or experts. Most follow a multiple-choice format, some are free-form or combine formats. (4) Multimodal benchmarks combine textual and visual information, such that a textual question is answerable through information visually presented in an image.
3.2 Analysis of Benchmark Papers
Figure 1 gives a schematic overview of our benchmark paper analysis procedure. Guided by our research questions and following a content analysis approach similar to Birhane et al. (2022), we firstly coded all of the benchmark papers, i.e., research articles or pre-prints introducing the benchmark (as the main contribution, as a byproduct to a technical work, or as a test split to a new corpus). Using MAXQDA (VERBI Software, 2024), our first author firstly highlighted sections relevant to our research questions and suggested preliminary annotation labels on the fly. After the first phase of annotations, labels were merged and categorized where possible to create a codebook. This initial codebook was discussed in a workshop with four participants, including two of the authors and two colleagues from the TU Berlin. Discussions regarding the codebook content later informed its refinement and finalization. [^5] During the workshop, the codebook was presented as a list of labels with short descriptions and the external participants were asked to annotate two benchmark papers by marking and labeling text spans using this list. It, however, required a long time for the new annotators to comprehend the list of possible labels and understand the type of insights we were looking for. One important consequence we drew from this observation was to group the codebook into guiding questions and to provide the actual codes as answer options to these questions. This helped to accelerate the on-boarding. The final codebook was reformatted and implemented as an online questionnaire via LimeSurvey (LimeSurvey Project Team / Carsten Schmitz, 2012), which we used for the second wave of annotations. [^6] The full questionnaire is available in our repository. Using the final coding schema, all 30 benchmark papers were re-annotated by our first author and one external annotator each. We distributed the co-annotation among 12 experts, of which nine were PhD students with a research focus on NLP and QA, two were Master’s students and one was a medical professional–all with a working knowledge of NLP and experience in reading scientific texts. All annotators (including workshop participants and respective authors) were aged between 25 and 60, originating from India, Pakistan, China, Germany, and Kazakhstan. Roughly one third identified as female. [^7] All annotators (including workshop participants) were informed about the conditions and rights (including GDPR) upon participating in our study and all provided their written consent prior to participation. Their demographic details were collected in a separate questionnaire including a separate informed consent form.
Figure 2: Quantitative data analysis process for the benchmark datasets.
3.3 Analysis of Benchmark Datasets
For the quantitative analysis of social bias within the benchmark datasets, we retrieved external information about entities (people, places, events, etc.) mentioned in the question-answer pairs from Wikidata, [^8] https://www.wikidata.org in particular, gender, occupation, religion, and location properties.
Our analyses comprise two different scenarios depicted in Figure 2: Scenario 1: The questions or answers of benchmarks like NaturalQuestions and TriviaQA include entities described in Wikipedia articles and respective identifiers (e.g., article titles or URLs) are provided. Using these identifiers, we queried the Wikipedia API [^9] https://www.mediawiki.org/wiki/API:Main_page to retrieve the corresponding Wikidata QIDs. Using SPARQL, [^10] https://www.w3.org/TR/rdf-sparql-query/ we then retrieved properties of interest for these QIDs directly from the Wikidata knowledge graph, e.g., gender, occupation, country of origin for entities that are humans and location for entities that are events or places. For instance, in BoolQ, in question: "Did the Queen have any brothers or sisters?", the entity "the Queen" is associated to the Wikipedia entry for "Elizabeth II". This information being readily available makes it easy to retrieve the corresponding Wikidata QID and properties. Scenario 2: For benchmarks that are not by default linked to Wikipedia, we applied Entity Linking (EL) against Wikidata using the ReFinED model (Ayoola et al., 2022). [^11] We used the implementation available here: https://github.com/amazon-science/ReFinED (license: Apache 2.0). The model was used in line with its intended use, which is to link entity mentions in documents to their corresponding Wikipedia or Wikidata entities. This allowed us to identify entities mentioned in the questions and retrieve their respective Wikidata identifiers. We then followed the same steps as in the first scenario. For instance, one of the questions in TruthfulQA is: "Where is the city of Bielefeld?" [^12] A correct answer to this question is ”Bielefeld is in Germany” and an expected incorrect answer is ”Bielefeld does not exist”. The entity linker identifies "Bielefeld" as a Wikidata entity with the QID "Q2112", allowing to query properties associated with it from the Wikidata knowledge graph (e.g., its coordinates). We excluded BioASQ-QA, MATH, and the multimodal benchmarks from the analysis, as identifying social biases within these benchmarks would necessitate additional domain-specific expertise or extensive annotation efforts beyond the scope of this study. Some other benchmark datasets were excluded if they yielded less than 30 results for each property. A total of 20 benchmark datasets were included in our final quantitative analysis.
As can be seen in Figure 4, encyclopedic and commonsense knowledge is most represented across all benchmarks (we summarize "everyday/world knowledge" under commonsense). Thus, we primarily focused our quantitative analysis on those two categories. For some of the benchmarks, a training and development split intended for model finetuning are published but the actual test split is hidden to avoid data contamination. In such cases, we analyzed the development split. Otherwise, we defaulted to the test split.
4 Results
4.1 Benchmark Paper Analysis Results
We obtained two sets of annotations for each of the 30 benchmark papers, one by an internal annotator (first author) and one by an external annotator (inter-annotator agreement: $\kappa$=.78; $SD$=.10). [^13] Cohen’s $\kappa$ was computed on the basis of all yes-no questions excluding the ”suggest other annotation” category.Throughout this section, we present the internal annotations unless otherwise specified and only discuss some of the differences between internal and external annotations (all external results are presented in Appendix B).
4.1.1 Benchmark Motivation
Increased task difficulty, novelty, and more realistic problems were the most frequent reported motivations behind the benchmarks. Other motivating factors mentioned were increased benchmark dataset size, explainability/interpretability, and domain-specificity. We found none of the benchmark papers to be motivated by better social representativeness. Note that the external annotator found SIQA to be aiming for social representativeness since it is framed as a social intelligence benchmark (Sap et al., 2019). However, we did not find any evidence that the intention was to improve representativeness in a demographic sense.
4.1.2 Benchmark Creation and Annotation
From the 30 analyzed benchmark papers, 20 of the benchmarks consist of human-authored items. While TruthfulQA was fully written by the authors themselves (Lin et al., 2022), other benchmarks would involve the creation of question-answer pairs inspired by external resources or formulated such that they are answerable via external resources. In 13 cases, some type of web source was used as a basis. Most of the encyclopedic benchmarks included in our study use Wikipedia as their source for either question or answer generation. SQuAD v1.1 (Rajpurkar et al., 2016) consists of more than 100,000 questions about Wikipedia articles, posed by crowdworkers. Similarly, for StrategyQA (Geva et al., 2021), HotpotQA (Yang et al., 2018), and TruthfulQA (Lin et al., 2022), crowdworkers created question-answer pairs inspired by Wikipedia content. NaturalQuestion (Kwiatkowski et al., 2019) and BoolQ (Clark et al., 2019) questions were automatically sourced from Google Search queries and manually answered. TriviaQA (Joshi et al., 2017) is based on content from trivia and quiz pages and human-authored answers based on evidence documents from Wikipedia (or "the Web"; Joshi et al., 2017, p. 1602). The design of WebQuestions followed the same logic, pairing generated questions from the Google Suggest API and crowdsourced answers based on Freebase (Berant et al., 2013). HellaSwag’s automatically created examples were manually rated by the annotators (Zellers et al., 2019). Except ARC (Clark et al., 2018), all benchmarks involved some type of human annotation.
4.1.3 Benchmark Language
All but one of the selected benchmarks were in English, only. However, only 12 of the benchmark papers explicitly state this information. In all other cases, we had to derive this information from data examples. For these cases, we have to assume that the recruited annotators were sufficiently capable of understanding and following English instructions and writing and labeling English data examples. An exception to English as a default, is XQuAD, a multilingual benchmark based on translations of the English SQuAD v1.1 (Spanish, German, Greek, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, and Hindi; Artetxe et al., 2020). Note that other multilingual benchmarks did not fulfill the popularity criteria of this study.
4.1.4 Annotator Recruitment Criteria
For 50% of the benchmarks, crowdworkers were hired through Amazon Mechanical Turk. [^14] https://www.mturk.com/ Other platforms used are Surge AI [^15] https://www.surgehq.ai/ (Cobbe et al., 2021) and Upwork [^16] https://www.upwork.com/ (Rein et al., 2023). Again, only 15 benchmark papers mention criteria for the selection of annotators (see Table 1). These would include performance on the task, e.g., appraised in a screening test (Reddy et al., 2019), or their ratings on the crowdworking platform (Rein et al., 2023). Sometimes annotators were recruited due to their availability as co-authors or colleagues (Gordon et al., 2012; Lin et al., 2022; Yue et al., 2024). Another reason for recruitment would be expertise in a certain domain. BioASQ-QA, for example, is a biomedical benchmark that is fully written by domain experts (Krithara et al., 2023). It is reported where and in what type of institutions the experts hold positions (European universities, hospitals, and research institutes) as well as their concrete areas of research (e.g., "cardiovascular endocrinology, psychiatry, psychophysiology, pharmacology", p. 3). In StrategyQA, the authors refer to themselves as expert annotators (Geva et al., 2021). In other instances, what defines an expert is less clear. For example, in the OpenBookQA benchmark paper it is stated that the data were "filtered by an in-house expert to ensure higher quality" (Mihaylov et al., 2018, p. 2384) without further elaboration.
Table 1: Annotator recruitment criteria and demographics. Abs. number of mentions across benchmark papers.
Criterion #
none 15
availability 3
task performance 6
domain expertise 4
other 3
Demographic #
none 17
country of origin 1
recruitment country 3
education 3
area of expertise 5
age 0
gender 0
ethnicity 0
other 2
4.1.5 Annotator Demographics
Out of the 29 benchmark papers involving human annotators, 17 failed to report any demographic information (see Table 1). Country of recruitment or origin was mentioned for SQuAD, DROP, OpenBookQA, and MATH, exclusively referring to the USA, Canada, or North America in general (Rajpurkar et al., 2016; Dua et al., 2019; Mihaylov et al., 2018; Hendrycks et al., 2021). Level of education was mentioned in OpenBookQA (Master’s; Mihaylov et al., 2018), GPQA (PhD or higher; Rein et al., 2023), and MMMU (college students; Yue et al., 2024), which are based on textbook problems or exam knowledge. Information on age, gender, and ethnicity were not identified in the benchmark papers (by internal nor external annotator). Another indicator for demographic aspects are the author affiliations. We found that those are centered around renown North-American research institutes, universities, and technology firms. In sum, 13 of the benchmark papers were co-authored by researchers affiliated to the Allen Institute for Artificial Intelligence (Allen AI) and 8 by researchers affiliated to the University of Washington (UW).
4.1.6 Benchmark Bias and Toxicity
We asked annotators to answer the following question and include evidence for their answer: "Are analyses of aspects related to social bias, representativeness or toxicity in the benchmark dataset reported and, if so, what type of analyses?" The external annotators identified 4 benchmarks as informative in this regard. However, we noticed that they appeared to work on a different understanding of bias than us. For instance, OK-VQA utilizes (non-specific) label balancing to avoid heuristic prediction behavior [^17] For example, the question ”What season is it?” was mostly accompanied by the answer ”Winter” incentivizing the model to default to this answer (Marino et al., 2019). and for NaturalQuestions, an in-depth analyses of annotation variability was conducted. This indeed can be done in a social bias-sensitive manner (Haliburton et al., 2024), but in this case the focus was on general annotation quality (Kwiatkowski et al., 2019). We, hence, count these cases as uninformative of social bias or toxicity aspects.
We finally identified 3 out of 30 benchmark papers that clearly flag social biases in their data. [^18] None of the benchmark papers mentioned any toxicity-related metric (full agreement between internal and external annotations). The WinoGender bias metric (Rudinger et al., 2018) was applied to models trained on the WinoGrande train split (Sakaguchi et al., 2021) to verify its relative gender-fairness. The QuAC datasheet mentions potential biases towards famous men in its dataset as well as other not further specified biases.191919quac.ai/datasheet.pdf The GPQA benchmark paper explicitly states that bias was not avoided during the dataset creation. The authors "make no claim that GPQA is a representative sample of any population of questions that are likely to come up in the course of scientific practice," (Rein et al., 2023, p. 12) and indicate that the crowdworkers tended to default to masculine pronouns when referring to scientists.
An additional keyword matching for the terms "diverse" and "diversity" yielded matches in two thirds of the benchmark papers: Several pay attention to domain or topic diversity (e.g., Geva et al., 2021; Lu et al., 2022; Lin et al., 2022), question or answer diversity (e.g., Zellers et al., 2019; Bisk et al., 2019; Artetxe et al., 2020), as well as lexical diversity (e.g., Reddy et al., 2019; Dua et al., 2019; Cobbe et al., 2021). Yet, again, none of them account for demographic diversity.
4.2 Benchmark Data Analysis Results
Next, we analyzed distributions of gender, occupation, religion and location properties found for entities across 20 benchmark datasets (see Table 2, Appendix A), following the procedure described in Section 3.2.202020The selection of demographic markers reflects dimensions that are frequently discussed in the social bias in NLP literature (Sheng et al., 2021). The absolute number of entities differs greatly between benchmarks (see Table 6, Appendix B) due to differences in dataset sizes or the nature of the contents, e.g., HotpotQA (>20k entities) is inherently related to Wikipedia and, thus, highly overlaps with Wikidata, but MMLU (<100 entities) is designed to include knowledge not found on Wikipedia [^21] For MMLU, we only matched entities of type class of anatomical entity, cell type, and the like. and the commonsense benchmark COPA (<100 entities) does mostly not rely on real-world entities in its examples. [^22] Example: ”The man dropped on the floor. What happened as a result? For other commonsense and scholarly benchmarks we were in fact able to retrieve large numbers of entities suitable for our analysis: E.g., 893 human and 365 fictional human entities in SIQA, 480 human and 99 fictional human instances in GSM8K, or 317 mentions of U.S. States in ScienceQA.
Figure 3: Gender ratio for entities in encyclopedic, commonsense, and scholarly QA & RC benchmarks.
4.2.1 Gender and Occupation
Figure 3 shows the male-versus-female gender proportions across benchmarks. We only included benchmark datasets for which we found more than 30 gender entries. Genders beyond the binary were none or close to none and, hence, also not illustrated in the plot. The most favorable gender ratios are found in WinoGrande and HellaSwag, which are both commonsense benchmarks. As discussed in Section 4.1.6, a low gender bias metric was reported by the WinoGrande authors and our results confirm this observation. All Wikipedia-based benchmarks, like DROP, SQuAD, or TriviaQA exhibit prominent gender gaps. In fact, the DROP benchmark dataset is only based on text passages about male-dominated "National Football League (NFL) game summaries and history articles" (Dua et al., 2019, p. 2371).
For CommonsenseQA, we only retrieved 28 male and 5 female entities, but we also ran a keyword matching on its question set and found 179 questions containing "he", "man" or "his" and only 49 containing "she", "woman", "her", or "hers". Examples are: "He was working hard on his sculpture, what was he practicing?" and "After she finished washing clothes, what did the woman do with them?" For questions where gender does not play a role for the task at hand, the dataset creators happened to default more to male subjects.
Additionally, we found that the most represented occupations differ for female and male entities. For example for GSM8K (school math problems) and WinoGrande (commonsense), the male top-10 occupations for both benchmarks include several athletic professions, while the female top-10 occupations are biased towards entertainment roles. Visualizations of occupation-related gender biases for eight different benchmarks are provided in Figure 5, Appendix C.
4.2.2 Religion and Location
As indicators of cultural and geographic context, we examined religions and locations. Here, we examined the benchmarks for which more than 30 religion properties were retrieved (ranging between 33 for BoolQ and 652 for TriviaQA). Christianity and instances of Christian religions rank highest across benchmarks. In fact, Christianity and/or Catholicism are among the top-3 religion labels for 14 out of 15 benchmarks. All distributions are visualized in Figure 6, Appendix C.
Across encyclopedic, commonsense, and scholarly benchmarks, most coordinates are located around North America and Western Europe and Eastern and Southern regions are less represented. For HotpotQA, TriviaQA, and NaturalQuestions slightly more coordinates are located on the South American, African, and Australian continents compared to the other benchmarks. Map visualizations for 17 benchmarks are provided in Figure 7, Appendix C. We also retrieved location names associated to entities in the datasets. Again, Western regions are more represented. E.g., for BoolQ and StrategyQA, the most frequently named location are the United States (56% and 31%) and the United Kingdom (9% and 15%), followed by Canada (2%) for BoolQ and Brazil, and Japan (4% each) for StrategyQA.
5 Discussion
Our study reveals significant gaps in transparency as well as harmful social biases across the 30 most popular QA and RC benchmarks. While two-thirds consist of human-authored examples and nearly all involve human annotation, half of them fail to disclose annotator demographics or recruitment criteria and all fail to report gender, age, or ethnicity. For the remaining cases, countries of origin or recruitment are predominantly North American, reflecting the Western institutional affiliations of the benchmark authors. While the QuAC paper stands out for its thorough reporting, others like MMLU (which is commonly referenced to market flagship models of famous tech firms [^23] https://openai.com/index/hello-gpt-4o/ [^24] https://www.anthropic.com/news/3-5-models-and-computer-use) lack all of the details we were looking for. The benchmarks overwhelmingly prioritize task difficulty or novelty aspects over social representativeness. The reliance on Wikipedia (known for its representational issues; Sun and Peng, 2021; Menking and Rosenberg, 2021; Tripodi, 2023) for encyclopedic benchmarks, perpetuates underrepresentation of marginalized communities and introduces biases in gender, occupation, geography, and religion. But also commonsense and scholarly benchmarks were found to default to male and Western examples. All but one benchmark are fully English, but less than half report this information in the paper, disregarding that they are indeed language-specific benchmarks (Bender, 2011).
Our findings highlight a systemic lack of diversity and transparency in widely used QA and RC benchmarks. These shortcomings perpetuate the development of technologies that produce harmful, discriminatory outcomes. Furthermore, our findings exemplify once more a "laissez-faire attitude" (Paullada et al., 2021, p. 4) prevalent in AI dataset creation, which must be addressed through the implementation of robust documentation, validation, and representation standards. While we acknowledge the growing discourse around better AI evaluation practices (Wallach et al., 2024), we emphasize that the conversation must prioritize social bias alongside validity and transparency.
6 Conclusion
Our work finds significant limitations regarding transparency and social representativeness in 30 popular QA and RC benchmarks. Many of these benchmarks lack information about annotator demographics, recruitment criteria, and language specificity. Many are, furthermore, biased in terms of gender, occupation, religion, and geographic representation. This has objectionable epistemological and ethical implications, e.g., by incentivizing the development of technologies that serve the needs of a privileged few. We highlight the need for rigorous documentation, validation, and representation standards in LLM benchmarking.
Limitations
Due to the lack of transparency across benchmarks, we were unable to investigate the causal relationship between the identity of those involved in the benchmark creation and the biases found in the QA and RC benchmark datasets through statistical testing. Due to the immense annotation efforts involved, our analyses were limited in scope. Future work should include a larger number and wider range of benchmarks to allow for more generalizable conclusions.
Acknowledgments
This work was mainly conducted during the first author’s Research Fellowship at Weizenbaum Institute, Berlin. This work was also supported by the German Research Foundation (DFG) project NFDI4DS under Grant No.: 460234259 and a NVIDIA Academic Hardware Grant.
Appendix A Full Benchmark Paper Checklist
Table 2 provides a full checklist regarding reported aspects, category, and inclusion in the dataset analysis across all benchmarks.
Table 2: Checklist of social bias-relevant aspects stated in the benchmark papers & inclusion in quant. analysis.
Domain Benchmark Language
Recruitment
criteria
Demographics
Social bias
or toxicity
Data
analysis?
Encyclopedic QuAC -
DROP -
XQuAD - -
SQuAD - -
HotpotQA - - -
StrategyQA - - -
COQA - - -
NaturalQuestions - - - -
TriviaQA - - - -
BoolQ - - - -
WebQuestions - - - -
Commonsense COPA - - -
WinoGrande -
PIQA - - - -
CommonsenseQA - - -
TruthfulQA - - -
HellaSwag - - -
SIQA - - - -
Scholarly BioASQ-QA - -
GPQA
RACE - -
OpenBookQA - -
MATH - - -
ScienceQA - - - -
MMLU - - - - -
GSM8K - - - -
ARC - - - -
Multimodal MMMU - -
TextVQA - - - - -
OK-VQA - - - - -
Appendix B Benchmark Paper Analysis Ext’d
Figure 4 provides an overview of the domain/ topic distribution across all benchmarks. Table 3 lists reported motivations across benchmarks and Table 4 the data sources. Table 5 shows the external annotations of annotator recruitment criteria and demographics (internal: Table 1).
Figure 4: Distribution of domains across benchmarks.
Table 3: Reported motivations. Abs. counts across papers. Internal (Int.) vs. external (Ext.) annotation.
Motivation Int. Ext.
increased difficulty 16 17
decreased difficulty 0 1
defining a new task 10 10
more realistic questions 9 10
better social representativeness 0 1
other 9 6
Table 4: Reported data sources. Abs. counts across papers. Internal (Int.) vs. external (Ext.) annotation.
Source Int. Ext.
human-authored 20 20
open access/ web data 13 14
reusing existing AI/NLP dataset 8 9
exams or textbooks 5 6
synthetic 1 1
proprietary/ internal source 0 0
other 1 2
Table 5: External annotations of annotator recruitment criteria and demographics. Abs. number of mentions.
Criterion #
none 14
availability 1
task performance 7
domain expertise 5
other 3
Demographic #
none 17
country of origin 1
recruitment country 2
education 4
area of expertise 3
age 1
gender 0
ethnicity 0
other 4
Appendix C Benchmark Dataset Analysis Ext’d
Table 6 lists detailed counts of entities extracted using the procedure described in Section 3.2. Figures 5 and 6 present relative frequencies of occupations by gender and religion [^25] Note that we replaced the term ”The Church of Jesus Christ of Latter-day Saints” with ”Mormon Church” for better proportions of the graph visualization. across benchmarks. Figure 7 illustrates the distributions of coordinates.
Table 6: Detailed list of the numbers of Wikidata entities and associated properties extracted for each benchmark.
#Entities #Extracted properties
Instance
of
Gender
Occu-
pation
Ethni-
city
Reli-
gion
Coordi-
nates
Location
names
Entity
linking?
Encycl. DROP 880 804 76 52 14 119 42 411 -
SQuAD 10570 9462 1173 1150 287 610 1242 4860 -
HotpotQA 22189 21077 6027 5684 103 541 3121 21103 -
StrategyQA 229 223 48 44 4 18 30 183
COQA 1349 1194 334 289 136 191 349 1264
NaturalQu. 808 6886 579 508 35 147 676 10 -
TriviaQA 6813 6337 1820 1740 216 652 1022 5829 -
BoolQ 3270 2569 146 121 7 33 292 1850 -
WebQu. 755 740 82 75 42 67 213 701 -
Comm. WinoGrande 799 774 477 356 26 63 56 809
Comm.QA 208 153 33 25 5 8 26 100
TruthfulQA 644 604 62 59 141 107 289 726
HellaSwag 3618 3228 309 270 201 106 417 2351
SIQA 2142 2132 1755 1531 148 147 76 2628
Schol. GPQA 310 274 16 17 13 3 28 99
RACE 1350 1215 411 370 147 145 349 1424
OpenB.QA 282 230 2 2 8 2 57 101
ScienceQA 2339 1820 453 346 56 101 554 1573
GSM8K 1096 1069 787 602 36 90 77 1239
ARC 695 570 54 44 18 12 111 338
Figure 5: Top-10 occupations by gender across benchmarks (if 300 or more occupations identified).
Figure 6: Top-10 religions found for entities across benchmarks (if 30 or more instances identified).
Figure 7: Distribution of coordinates found for entities across benchmarks (if 30 or more instances identified).